Endicott Enterprises, Inc. has issued thirty-year semiannual coupon bonds with a face value of $1,000. If the annual coupon rate is 14% and the current yield to maturity is 8%, what is the firm’s current price per bond?

Please use a calculator for a full explication of FV, PMT, N , I/Y

Thank you
Answer #1

FV = Future value = 1000

PMT = Annual payment = 1000*14%/2 = 70

N = Number of periods = 30*2 = 60

I/Y = Yield to maturity = 8%/2 = 4%

PV = present value = ?

Answer: 1678.70

---------------------------------------------------